OpenStudy (rockycatsrule):

What is the solution to the system of equations? 5x+4y=8 x-2y=10 (_______ ,________) plz help me plz

OpenStudy (error1603):

you have to equal x or y to solve so, 5x+4y=8 x-2y+10 (2) 5x+4y=8 2x-4y=20 ======= 7x=28 x=28/7 x=4

OpenStudy (rockycatsrule):

i am confused so is the answer 28 /7 or 4/7???

OpenStudy (error1603):

+4y and -4y will be cancelled.

OpenStudy (rockycatsrule):

k

OpenStudy (error1603):

to find y,you use any one original equation and plug in the value of x 5(4)+4y=8 20+4y=8 4y=8-20 4y=-12 y=-12/4 y=-3

OpenStudy (error1603):

Understand?

OpenStudy (rockycatsrule):

no

OpenStudy (error1603):

x=4.y=-3

OpenStudy (error1603):

@rockycatsrule